OpenStudy (anonymous):

2. What is the value of 4C2? (1 point) 6 8 16 80

OpenStudy (thewafflebro):

potato

OpenStudy (welshfella):

4! ----- 2! 2!

OpenStudy (anonymous):

@TheWaffleBRO can u leave me alone?!!

OpenStudy (thewafflebro):

idk

OpenStudy (anonymous):

@welshfella whats the answer?

OpenStudy (thewafflebro):

byee

OpenStudy (welshfella):

a quick way to do these combinations is 4 *3 --- 2 * 1

OpenStudy (anonymous):

wait im doing the problem

OpenStudy (welshfella):

you can work out 4*3 / 2*1 i'm sure

OpenStudy (welshfella):

12/2 = ?

OpenStudy (anonymous):

12/2 ?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

so it would be 6 right?

rvc (rvc):

\[\left(\begin{matrix}4 \\ 2\end{matrix}\right)=\frac{ 4! }{ (4-2)!2!}\]

OpenStudy (welshfella):

right

OpenStudy (anonymous):

thanks :)

OpenStudy (welshfella):

the quick way is like this